# Large OCR Model: An Empirical Study of Scaling Law for OCR

### 02 Jan 2024 v2

@@ 1 Introduction
Optical Character Recognition (OCR) is a technology designed to detect and interpret textual content within images, such as scanned documents and photographs. The field of OCR has undergone significant evolution in recent years, largely driven by advancements in deep learning. These advancements have enabled OCR system to be successfully deployed in a variety of real-world applications, including the automatic recognition of license plates, official documents and bills. A typical OCR system operates in two primary phases: text detection and text recognition. In this context, our focus is on the text recognition phase, which involves identifying and extracting text from predefined bounding boxes.
@@ 3.1 Model Scaling
TrOCR [42] is an end-to-end OCR model that utilizes the Transformer architecture. It integrates pre-trained Computer Vision (CV) and NLP models. And it is the first work that jointly leverages image Transformer and text Transformer for text recognition tasks in OCR. The scaling laws of Transformer language models [31] and Vision Transformers [94] have been studied, the scaling laws for models in the OCR field have not yet been explored. Based on this, we scaled the TrOCR model size and attempted to analyze the accuracy change curves of models with varying sizes.
@@
PARSeq [8] follows an encoder-decoder architecture. The encoder part utilizes the Vision Transformer (ViT) model to extract image features, while the decoder follows the same architecture as the pre-LayerNorm [83]. Transformer decoder in this study utilizes twice the number of attention heads, where nhead = dmodel/32. In contrast to the standard ViT, the encoder removes the token and inputs all the output tokens into the decoder.
@@ 3.2 Dataset
To augment the diversity of the generated corpus, we sourced 700,000 corpus entries from the most extensively utilized English corpus website globally333https://www.english-corpora.org/corpora.asp. For the background selection in our synthesized images, we employed natural scene pictures provided by SynthText as the backdrop. Utilizing these two synthesis methods, we successfully synthesized a total of 60M data samples.
@@ 4.1 Smooth Power Laws
Scaling TrOCR Models We trained 4 different scales (ranging in size from 43.09M to 1B) of TrOCR models on MJ and ST. The experimental results on 6 common test benchmarks are shown in Table 4. As shown in Fig 1(a), our analysis reveals a linear relationship on the log-log plot between the parameter count and modeling performance. This relationship can be described by a power-law equation (). Employing Algorithm 1, we utilized the first three models (TrOCR-S, TrOCR-B and TrOCR-L) to obtain the power function equation . The last model (TrOCR-H) accurately aligns with the fitted straight line, demonstrating the effectiveness of the power law. The power law of the TrOCR model is as follows.
@@ 4.2 Other Observations
Large-scale models make better use of samples. As we continue to enlarge the model size, the model accuracy improves consistently. The phenomenon can be observed in Table 4 and 5. To improve training results, we can modify recognition models built on the Transformer architecture by utilizing the scaling laws of the vision Transformer. As shown in Figure 4 with respect to the total number of images “seen” during PARSeq training stage of different sizes, it is clear that larger models utilize samples more effectively than their smaller models. When PARSeq models of different sizes are trained with an equal number of samples, smaller models exhibit a higher error rate compared to larger models. Furthermore, we observed that larger models tend to require fewer epochs to converge. For instance, PARSeq-S reached optimal accuracy in 32 epochs, whereas PARSeq-B needed only 14 epochs, and PARSeq-L just 5 epochs. These findings suggest that with adequate training resources, it is more beneficial to train a larger model for fewer steps. This mirrors similar findings in language modelling [31] and machine translation [22]. However, when training time is a limiting factor, opting for a smaller model may be more practical.
The proportion of training data from various sources is crucial for model training. The REBU-Syn comprises both real and synthetic data. According to prior studies [8, 96], real data typically outperforms synthetic data in training efficiency, though synthetic data still plays a valuable role. Due to the high costs associated with obtaining and labeling real data, which often do not meet the volume requirements for model training, reliance on synthetic data is necessary. However, the effectiveness of synthetic data raises a question: Does more synthetic data always equate to better performance? Our findings suggest that an optimal ratio between real and synthetic data is crucial for enhanced model performance.
@@
Task-related pre-trained models are more effective. The utility of pretrained models in low-level vision tasks is well-known, but their applicability in OCR tasks warrants investigation. To address this, we experimented with various pretrained models, some trained on ImageNet and others specifically for OCR recognition tasks. In the last two rows of Table 10, we maintained consistent training schedules, learning rates and epochs as used for PARSeq. Intriguingly, the ImageNet-21k pre-trained models underperformed compared to those trained from scratch, a trend observed in both PARSeq and CLIP4STR models. This suggests that pretraining on non-OCR-specific tasks might not be beneficial, and can even be detrimental to OCR performance. The OCR task necessitates a connection between visual and textual elements, akin to the CLIP experiment’s original purpose, whereas purely visual tasks focus more on high-level semantics and lack the textual nuances critical for OCR.
@@ Appendix D OCR Enhanced LMM
In the realm of large-scale models, we observe a distinct bifurcation into two primary categories: Large Language Models (LLMs) and Large Multimodal Models(LMMs). It is crucial to acknowledge that while LLMs are devoid of a visual component, LMMs’ visual branches demonstrate room for enhancement in terms of OCR capabilities[73]. This observation underscores the relative underdevelopment of OCR proficiency within large-scale models. However, OCR tasks offer a promising avenue to address this shortfall, thereby motivating our investigation into the benefits of integrating OCR with these models.
@@
Result and Analyze We performed a detailed comparative analysis to assess the accuracy of the QWen-VL-chat model, examining its performance with and without OCR integration, as delineated in Table 5. Our results reveal a significant improvement in the accuracy of the model for scene-based VQA tasks upon the integration of OCR. Additionally, there is a noticeable enhancement in document-based VQA tasks. These findings suggest that the incorporation of OCR not only enhances the model’s accuracy but also extends its generalization capabilities across diverse VQA scenarios. This evidence distinctly highlights the vital role that OCR inputs play in augmenting the performance of LMM for downstream tasks. Furthermore, the improved accuracy with OCR integration underscores the model’s enhanced ability to interpret and analyze combined visual and textual data, thereby validating the efficacy of multimodal approaches in tackling complex analytical challenges.
VQA Visulization Analysis Our visual analysis of QWen-VL-Chat, with and without the OCR module, across varied datasets offers critical insights. In natural scene VQA tasks, QWen-VL-Chat encounters difficulties in detecting small text in images. The upper left corner of Fig 2, the model overlooks pertinent content, erroneously indicating its absence. Moreover, its tendency to inaccurately complete blurred text stems from its sophisticated linguistic abilities. This is evident in the upper right corner of Fig 2, where ’dr’ in ’45th parallel dr’ is incorrectly expanded to ’drive’. Notably, the model’s text misidentification issues, such as converting ’honghe’ to ’Hongte’ on a cigarette pack as depicted in the lower left corner of Fig 2 (mistaking the second ’h’ for a ’t’), are significantly mitigated by integrating our OCR module.

### 31 Jan 2024 v3

# An Empirical Study of Scaling Law for OCR

@@ 1 Introduction
Optical Character Recognition (OCR) is a technology designed to detect and interpret textual content within images, such as scanned documents and photographs. A typical OCR system operates in two primary phases: text detection and text recognition. In this context, our focus is on the text recognition phase, which involves identifying and extracting text from predefined bounding boxes. In text recognition tasks, it is subdivided into scene text recognition and scanned document recognition. Scene text recognition is more challenging than traditional scanned documents because it needs to solve more complex problems in practical applications, such as illumination changes, occlusion, distortion, angle changes and other factors. The impact makes the identification task more difficult. Moreover, STR, as an active research field, has greater room for technological improvement and innovation. Consequently, in order to investigate the applicability of the scaling law, this paper focuses on the scene text recognition task.
@@ 3.1 Model Scaling
TrOCR [43] is a text recognition model that utilizes pure Transformer architecture. It integrates pre-trained Computer Vision (CV) and NLP models. And it is the first work that jointly leverages image Transformer and text Transformer for text recognition tasks in OCR. The scaling laws of Transformer language models [32] and Vision Transformers [96] have been studied, the scaling laws for models in the OCR field have not yet been explored. Based on this, we scaled the TrOCR model size and attempted to analyze the accuracy change curves of models with varying sizes.
@@
PARSeq [8] follows an encoder-decoder architecture. PARSeq is also based on the Transformer framework with excellent accuracy, which perfectly fits the scope of our research. The encoder part utilizes the Vision Transformer (ViT) model to extract image features, while the decoder follows the same architecture as the pre-LayerNorm [84]. Transformer decoder in this study utilizes twice the number of attention heads, where nhead = dmodel/32. In contrast to the standard ViT, the encoder removes the token and inputs all the output tokens into the decoder.
@@ 3.2 Dataset
To augment the diversity of the generated corpus, we sourced 700,000 corpus entries from the most extensively utilized English corpus website globally333https://www.english-corpora.org/corpora.asp. For the background selection in our synthesized images, we employed natural scene pictures provided by SynthText as the backdrop. Utilizing these two synthesis methods, we successfully synthesized a total of 60M data samples. Code for data synthesis is available444https://github.com/large-ocr-model/large-ocr-model.github.io.
@@ 4.1 Smooth Power Laws
Scaling TrOCR Models We trained 4 different scales (ranging in size from 43.09M to 1B) of TrOCR models. In order to maintain fairness and consistency with the experimental setting in the original TrOCR paper, we use MJ and ST to train TrOCR models with different model sizes. The experimental results on 6 common test benchmarks are shown in Table 4. As shown in Fig 1(a), our analysis reveals a linear relationship on the log-log plot between the parameter count and modeling performance. This relationship can be described by a power-law equation (). Employing Algorithm 1 in the appendix, we utilized the first three models (TrOCR-S, TrOCR-B and TrOCR-L) to obtain the power function equation . The last model (TrOCR-H) accurately aligns with the fitted straight line, demonstrating the effectiveness of the power law. The power law of the TrOCR model is as follows.
@@ 4.2 Other Observations
Large-scale models make better use of samples.
As we continue to enlarge the model size, the model accuracy improves consistently. The phenomenon can be observed in Table 4 and 5. To improve training results, we can modify recognition models built on the Transformer architecture by utilizing the scaling laws of the vision Transformer. As shown in Figure 4 with respect to the total number of images “seen” during PARSeq training stage of different sizes, it is clear that larger models utilize samples more effectively than their smaller models. When PARSeq models of different sizes are trained with an equal number of samples, smaller models exhibit a higher error rate compared to larger models. Furthermore, we observed that larger models tend to require fewer epochs to converge. For instance, PARSeq-S reached optimal accuracy in 32 epochs, whereas PARSeq-B needed only 14 epochs, and PARSeq-L just 5 epochs. These findings suggest that with adequate training resources, it is more beneficial to train a larger model for fewer steps. This mirrors similar findings in language modelling [32] and machine translation [22]. However, when training time is a limiting factor, opting for a smaller model may be more practical.
The proportion of training data from various sources is crucial for model training.
The REBU-Syn comprises both real and synthetic data. According to prior studies [8, 97], real data typically outperforms synthetic data in training efficiency, though synthetic data still plays a valuable role. Due to the high costs associated with obtaining and labeling real data, which often do not meet the volume requirements for model training, reliance on synthetic data is necessary. However, the effectiveness of synthetic data raises a question: Does more synthetic data always equate to better performance? Our findings suggest that an optimal ratio between real and synthetic data is crucial for enhanced model performance.
@@
Task-related pre-trained models are more effective.
The utility of pretrained models in low-level vision tasks is well-known, but their applicability in OCR tasks warrants investigation. To address this, we experimented with various pretrained models, some trained on ImageNet and others specifically for OCR recognition tasks. In the last two rows of Table 10, we maintained consistent training schedules, learning rates and epochs as used for PARSeq. Intriguingly, the ImageNet-21k pre-trained models underperformed compared to those trained from scratch, a trend observed in both PARSeq and CLIP4STR models. This suggests that pretraining on non-OCR-specific tasks might not be beneficial, and can even be detrimental to OCR performance. The OCR task necessitates a connection between visual and textual elements, akin to the CLIP experiment’s original purpose, whereas purely visual tasks focus more on high-level semantics and lack the textual nuances critical for OCR.
@@ 5 Discussion and Conclusion
Our experiments are based on large-scale natural scene text data sets. In the future, we will consider exploring the scaling law in more challenging text recognition data sets such as handwriting and historical texts.
@@ Appendix D OCR Enhanced LMM
In the realm of large-scale models, we observe a distinct bifurcation into two primary categories: Large Language Models (LLMs) and Large Multimodal Models(LMMs). It is crucial to acknowledge that while LLMs are devoid of a visual component, LMMs’ visual branches demonstrate room for enhancement in terms of OCR capabilities[74]. This observation underscores the relative underdevelopment of OCR proficiency within large-scale models. Optical Character Recognition (OCR) tasks, however, offer a promising avenue to address this shortfall, thereby motivating our investigation into the benefits of integrating OCR with these models.
@@
Result and Analyze We performed a detailed comparative analysis to assess the accuracy of the QWen-VL-chat model, examining its performance with and without OCR integration, as delineated in Table 5. Our results reveal a significant improvement in the accuracy of the model for scene-based VQA tasks upon the integration of OCR. Additionally, there is a noticeable enhancement in document-based VQA tasks. These findings suggest that the incorporation of OCR not only enhances the model’s accuracy but also extends its generalization capabilities across diverse VQA scenarios. This evidence distinctly highlights the vital role that OCR inputs play in augmenting the performance of LVLM for downstream tasks. Furthermore, the improved accuracy with OCR integration underscores the model’s enhanced ability to interpret and analyze combined visual and textual data, thereby validating the efficacy of multimodal approaches in tackling complex analytical challenges.
VQA Visulization Analysis Our visual analysis of QWen-VL-Chat, with and without the OCR module, across varied datasets offers critical insights. In natural scene Visual Question Answering (VQA) tasks, QWen-VL-Chat encounters difficulties in detecting small text in images. The upper left corner of Fig 2, the model overlooks pertinent content, erroneously indicating its absence. Moreover, its tendency to inaccurately complete blurred text stems from its sophisticated linguistic abilities. This is evident in the upper right corner of Fig 2, where ’dr’ in ’45th parallel dr’ is incorrectly expanded to ’drive’. Notably, the model’s text misidentification issues, such as converting ’honghe’ to ’Hongte’ on a cigarette pack as depicted in the lower left corner of Fig 2 (mistaking the second ’h’ for a ’t’), are significantly mitigated by integrating our OCR module.
@@
Appendix E Scaling law algorithm description
We formalize the power law of performance in terms of scaling factors, and the implemented details are shown in Algorithm 1.
Appendix F The scaling law on Union14M benchmark
We supplement the experiments with scaling laws on the Union14M benchmark. The parameters and accuracy of PARSeq-(S/B/L) and CLIP4STR-(S/B/L) on the Union14M benchmark are shown in Table 7 and Table 6 respectively. The curves of scaling law on CLIP4STR and PARSeq models are shown in Fig 4. It demonstrates that the scaling law is still applicable on the union14M benchmark.
| Method | Param (M) | Avg |
| PARSeq-S | 22.5 | 89.89 |
| PARSeq-B | 104.0 | 90.37 |
| PARSeq-L | 335.9 | 90.81 |
| Method | Param (M) | Avg |
| CLIP4STR-S | 43.6 | 91.90 |
| CLIP4STR-B | 86.7 | 92.08 |
| CLIP4STR-L | 268.2 | 92.19 |
Appendix G Applicability in document contexts
We validate the power law using scaling model sizes on the document dataset in addition to the STR benchmark. The FUNSD [30] dataset contains a large number of scanned documents, and each sample is annotated with detailed text, word bounding boxes, and structured tags. It is intended to support the development and assessment of model performance by researchers for the purpose of processing and comprehending information from scanned documents in noisy, real-world. Notably, CLIP4STR-L* achieved a SOTA accuracy of 96.5, surpassing the previous best, CLIP4STR-L. The experimental results are shown in Table 8. These results highlight the robustness of CLIP4STR-L* in both scene and document text recognition tasks.
| Model | Word Acc |
| CLIP4STR-L | 96.02 |
| CLIP4STR-L* | 96.50 |